OpenStudy (anonymous):

find 3 consecutive positive even integers such that the product of the second and third is 80

OpenStudy (anonymous):

let x-2,x,x+2 be consecutive even integers x(x+2)=80 x^2+2x-80=0 make factors and find the value of x

OpenStudy (anonymous):

how about this one? find three consecutive positive integers such that the square of the smallest exceeds the largest by 10 I really appreciate your help

OpenStudy (anonymous):

let x-1,x,x+1 be consecutive integers and x-1 be smallest,then (x+1)^2-(x-1)^2=10 solve it